OpenStudy (anonymous):

Joshua is creating a series of rectangles, one inside the other. The length of the first rectangle, Rectangle A, is five inches. Joshua dilates Rectangle A about its center by a scale factor of two to get Rectangle B. He repeats this process using the same scale factor till he gets a rectangle of length 80 inches. How many dilations did Joshua perform including the dilation of Rectangle A? 4 2 8 16

OpenStudy (anonymous):

his original length is 5 1st dilation: 5*2 = 10 2nd dilation: 5*2*2 = 20 3rd dilation: 5*2*2*2 = 40 4th dilation: 5*2*2*2*2 = 80

OpenStudy (anonymous):

o_o . is it that yur super duper smart or my brain is not working today

OpenStudy (anonymous):

no.... the above is just a guess

OpenStudy (anonymous):

yea but it makes sense.

OpenStudy (anonymous):

my guesses are good but when I actually do the work, it's wrong...

OpenStudy (anonymous):

i'm kidding!!! :)

OpenStudy (anonymous):

no i did th work its right.

OpenStudy (anonymous):

lol can you help me with 2 more please?

OpenStudy (anonymous):

that means you're smart...

OpenStudy (anonymous):

post it up... we'll help you... :)

OpenStudy (anonymous):

i knoe my moms tells me all the time its js my mind is everywhere today i usually dnt need this much help .

OpenStudy (anonymous):

and thank you very much ill post the next one now

OpenStudy (anonymous):

see? mother knows best...